### v1

# A Comprehensive Characterization of Galaxy-Cool CGM Connections at $z<0.4$ with DESI Year 1 Data

Abstract
We investigate the relationships between the properties of the cool circumgalactic medium (CGM), traced by Ca II absorption lines, and those of galaxies at $z<0.4$ by utilizing a galaxy-quasar pair sample compiled from the Year 1 data of the Dark Energy Spectroscopic Instrument (DESI). This large dataset, containing $\sim 900,000$ galaxy-quasar pairs within $200\,\rm kpc$, enables us to obtain composite spectra with sensitivity reaching to $\rm m\AA$ level and to explore the Ca II absorption as a function of stellar mass, SFR, redshift, and galaxy types, including AGNs. Our results show a positive correlation between the absorption strength and stellar mass of star-forming galaxies with $\langle W_{0}^{\rm Ca\ II}\rangle\propto M_{*}^{0.5}$ over three orders of magnitude in stellar mass from $\sim 10^{8}$ to $10^{11}\,M_{\odot}$, while such a mass dependence is weaker for quiescent galaxies. For galaxies with similar mass, we find that Ca II absorption is stronger around star-forming galaxies than around quiescent galaxies especially within the inner regions ($<30\,\rm kpc$) of the halos. Among star-forming galaxies, the Ca II absorption further correlates with SFR, following $\propto\mathrm{SFR^{0.3}}$. However, in contrast to the results at higher redshifts, we find that stronger absorption around star-forming galaxies is not preferentially observed along the minor axis of galaxies, indicating a possible redshift evolution of CGM dynamics resulting from galactic feedback. Moreover, no significant difference between the properties of the cool gas around AGNs and galaxies is detected. Finally, we measure the absorption profiles with respect to the virial radius of dark matter halos and estimate the total Ca II mass in the CGM. The results show that the CGM contains a metal mass comparable to the metal mass in the ISM of galaxies.
Circumgalactic medium (1879), Spectroscopy (1558), Extragalactic astronomy (506)
1 Introduction
The circumgalactic medium (CGM) contains signatures of complex physical processes that regulate gas flowing in and out of galaxies and drive galaxy evolution (see Tumlinson et al., 2017; Péroux & Howk, 2020; Faucher-Giguère & Oh, 2023, for reviews on the topic of the CGM). Therefore, observed properties of the CGM and its connection to the properties of galaxies have been considered as crucial constraints on the models of galaxy evolution, especially the subgrid models implemented for the feedback processes driven by the explosions of massive stars and the activities of supermassive black holes (e.g., Liang et al., 2016; Ford et al., 2016; Anglés-Alcázar et al., 2017; Oppenheimer et al., 2018). To probe the properties of the CGM, absorption line spectroscopy has been widely used (e.g., Bahcall & Spitzer, 1969; Bergeron, 1986; Chen et al., 2010; Nielsen et al., 2013; Tumlinson et al., 2013; Zhu et al., 2014; Lan et al., 2014; Anand et al., 2021; Bouché et al., 2025). By observing the spectra of background sources intercepting the CGM of galaxies, one can detect absorption line features induced by various transitions of elements, and investigate the properties of the circumgalactic gas, such as the radial distribution, kinematics, abundance, metallicity, and gas density, and explore how they correlate with the properties of the galaxies (e.g., Zych et al., 2009; Zhu et al., 2014; Lan et al., 2014; Nielsen et al., 2017; Pointon et al., 2019; Lan, 2020; Anand et al., 2021; Nateghi et al., 2024; Lehner et al., 2024; Cherrey et al., 2025).
Multi-wavelength observations are required to investigate the galaxy-CGM connections at different redshifts owing to the accessibility of strong absorption-line transitions. At $z<0.3$, UV spectrographs, COS and STIS (Green et al., 2012; Woodgate et al., 1998), on the Hubble Space Telescope have been used to detect absorption species, such as H I, O I, O VI, C II, C IV, Si II, and Si III, tracing gas in various physical conditions (e.g., Tumlinson et al., 2013; Werk et al., 2013, 2014; Bordoloi et al., 2014; Liang & Chen, 2014; Borthakur et al., 2015, 2016; Werk et al., 2016; Prochaska et al., 2017; Chen et al., 2020, 2020; Tchernyshyov et al., 2023; Zheng et al., 2024). These measurements have led to new explorations of the nature and origins of the multiphase CGM. However, given the sensitivity of the current instruments, only bright quasars can yield UV spectra with sufficient sensitivity for absorption line studies (e.g., Tumlinson et al., 2013). This in turn limits the number of galaxy-quasar pairs and the parameter space of galaxy properties available for galaxy-CGM correlation analysis. On the other hand, at $z\gtrsim 0.4$, optical spectra can be used to investigate the connection between galaxies and the cool CGM traced by Mg II absorption lines, one of the key accessible absorption species at optical wavelengths (e.g., Nestor et al., 2005; Chen et al., 2010; Nielsen et al., 2013; Zhu & Ménard, 2013a; Anand et al., 2021; Huang et al., 2021; Chen et al., 2023; Guha et al., 2024; Zou et al., 2024; Bouché et al., 2025). With large optical spectroscopic datasets provided by sky surveys, such as the Sloan Digital Sky Surveys (SDSS; York et al. 2000) and the Dark Energy Spectroscopic Instrument (DESI; Levi et al. 2013), statistical samples, consisting of thousands of galaxy-quasar pairs, have been built and used to measure the properties of the cool CGM as a function of galaxy properties, including stellar mass, star-formation rate (SFR) and AGN activity, across redshifts (e.g., Zhu & Ménard, 2013a; Lan et al., 2014; Lan, 2020; Anand et al., 2021; Chang et al., 2024; Wu et al., 2024), revealing correlations between the cool CGM and galaxy properties. However, observational measurements of these correlations at low redshifts are still limited, and therefore whether the relationships between galaxies and the cool CGM detected at $z\gtrsim 0.4$ exist at lower redshifts remains to be addressed.
[…]
In this work, we utilize the large spectroscopic data provided by the DESI survey (DESI Collaboration et al., 2016a, b) to measure the properties of the cool CGM traced by Ca II as a function of galaxy properties, including SFR, stellar mass, redshift, and azimuthal angle, covering a parameter space that has rarely been probed previously. We describe our sample and the analysis procedure to obtain composite spectra in Section 2. In Section 3, we present our results. In Section 4, we compare our results with previous studies and discuss the implications of our findings. Finally, we conclude our results in Section 5. Throughout this paper, we adopt a flat $\Lambda$CDM model with $h=0.7$ and $\Omega_{m}=0.3$.
2 Data analysis
2.1 DESI galaxy and quasar catalogs
The Dark Energy Spectroscopic Instrument (DESI) project is designed to measure the expansion rate of the Universe across cosmic time by using galaxies, quasars, and gas as tracers of large-scale structure at different redshifts (Levi et al., 2013; DESI Collaboration et al., 2022). To this end, a dedicated spectroscopic observational program has been developed utilizing an instrument consisting of 10 spectrographs and 5000 fibers (Silber et al., 2023; Miller et al., 2024). DESI primarily obtains spectra of four types of extragalactic sources at different redshift ranges, including galaxies that can be observed during the bright time targeted by Bright Galaxy Survey ($z<0.6$; Hahn et al. 2023), luminous red galaxies ($1.00.4<z<$ 1.0; Zhou et al. 2023), emission-line galaxies ($1.60.6<z<$ 1.6; Raichoor et al. 2023) and quasars ($z<3$; Chaussidon et al. 2023). These sources are selected based on the images of the DESI Legacy Imaging Surveys (Dey et al., 2019). The wavelength coverage of DESI spectra is from 3600 to 9800 Å with the spectral resolution varying from $R\sim 2000$ to 5000 for the blue camera to the NIR camera. The raw spectroscopic data are processed with a pipeline developed by Guy et al. (2023). In addition, an automatic pipeline, called redrock [^1] https://github.com/desihub/redrock, is developed and used for determining the redshifts of the sources from the spectra (Bailey et al. 2024; see also Anand et al. 2024). In order to maximize the efficiency of the survey, dedicated algorithms and codes have also been developed for target selection (Myers et al., 2023), fiber assignments (Raichoor et al., 2024) and survey operations (Schlafly et al., 2023). Prior to the main spectroscopy survey, DESI has undergone the “Survey Validation (SV)” (DESI Collaboration et al., 2024a) to validate the performance of the target selections. The five-month spectroscopic data from Survey Validation was released in June 2023 as part of Early Data Release (DESI Collaboration et al., 2024b).
Figure 1: SFR-mass distribution of the foreground galaxies. The solid line divides the samples into star-forming and quiescent galaxies. The contours mark the levels of numbers at 10, 100, 500, 1000, 5000, 8000.
In this work, we use the Year 1 dataset of the DESI survey (DESI Collaboration et al., 2024), which has yielded new measurements for constraining the expansion history of the Universe from $z=0$ to 2 (DESI Collaboration et al., 2024c, d, e, f, g, h; DESI Collaboration et al., 2025). More specifically, we construct our foreground galaxy and background source samples with the following two components:
  • Bright Galaxy Survey (BGS): We use galaxies with $0.40.02<z_{\mathrm{BGS}}<$ 0.4 observed in the bright galaxy survey as our foreground galaxies. The BGS sample consists of two selections, BGS bright and BGS faint samples. The BGS bright sample is selected with $r<19.5$ mag, and the BGS faint sample is selected with magnitude and color cuts (Hahn et al., 2023). Moreover, following the criteria from Hahn et al. (2023) and Lan et al. (2023), we select galaxies with $\Delta\chi^{2}\geq 40$ (the difference of $\chi^{2}$ between the second and the first best-fit spectral models from redrock).
    Besides the redshift information, we use the physical properties of the BGS galaxies, including the stellar mass ($M_{*}$) and star-formation rate (SFR), provided by Siudek et al. (2024). The authors use the CIGALE package (Boquien et al., 2019) for spectral energy distribution fits to observed magnitudes in the $g$, $r$, $z$, WISE 1 and WISE 2 bands extracted by the Tractor algorithm (Lang et al., 2016). The final fitting results exclude sources with low signal-to-noise ratio in photometry and poor fitting quality. For details, we refer the readers to Siudek et al. (2024).
    Figure 1 shows the stellar mass and SFR distribution of the BGS galaxy sample used in this analysis. To separate star-forming galaxies and quiescent galaxies, we adopt the following relation (black solid line) from Zhu & Ménard (2013b), which was derived from a sample of galaxies with a mean redshift $z\sim 0.1$,
    $\log_{10}{\frac{\mathrm{SFR}}{M_{\odot}\,\mathrm{yr^{-1}}}}=-0.79+0.65\,\left(\log_{10}{\frac{M_{*}}{M_{\odot}}}-10.0\right).$ (1)
    As shown in the figure, the BGS sample includes star-forming galaxies with $M_{*}$ down to $10^{7}-10^{8}\,M_{\odot}$ allowing us to probe the CGM of galaxies with mass across 3-4 orders of magnitude. The median mass uncertainty of the sample is 0.12 dex. For SFR, the typical uncertainties are 0.3 dex for star-forming galaxies and 8.1 dex for quiescent galaxies. We note that the reported uncertainty is in logarithmic scale. For quiescent galaxies with low estimated SFR ($\log_{10}\mathrm{SFR}/M_{\odot}\rm yr^{-1}\sim-7$), the corresponding uncertainty in linear scale is $\sim 10^{-6}M_{\odot}/\rm yr$. This precision is sufficient for classifying galaxies into star-forming and quiescent ones.
    In addition, we construct an AGN sample by selecting sources located in the AGN and composite regions (Kewley et al., 2006) of the BPT diagram (Baldwin et al., 1981). For this selection, we require a signal-to-noise ratio greater than 2 for the $\rm H\alpha$, $\rm H\beta$, [O III] 5007, and [N II] 6584 emission lines. The flux measurements for these lines are provided by the FastSpecFit [^2] https://fastspecfit.readthedocs.io/en/latest/ Spectral Synthesis and Emission-Line Catalog (FastSpecFit 2.1) (Moustakas et al., 2023).
  • Quasars: For the background sources, we use the quasar sample which is selected with a random forest algorithm combining magnitude and color information (Chaussidon et al., 2023). Similar to the BGS sample selection, we apply $\Delta\chi^{2}\geq 20$ to obtain sources with robust redrock redshift measurements (Alexander et al., 2023).
With the galaxy and quasar samples, we construct the galaxy-quasar pair sample by requiring (1) $z_{\mathrm{QSO}}-z_{\mathrm{BGS}}>0.1$ and (2) $\left(z_{\mathrm{BGS}}+1\right)\times 3800\,\mathrm{\AA}>\left(z_{\mathrm{QSO}}+1\right)\times 1216\,\mathrm{\AA}$, ensuring that the Ca II lines fall outside the Lyman alpha forest of the background quasar to avoid contamination. This yields a sample with impact parameter $r_{p}\leq 200\,\mathrm{kpc}$ having 855,703 galaxy-quasar pairs with 623,053 unique BGS galaxies and 496,756 unique QSOs. The median redshift of the unique BGS galaxies is $\langle z\rangle=0.1712$.
Figure 2: Examples of composite spectra as a function of impact parameters. The shades underneath are the associated uncertainties via bootstrapping. The solid blue lines are the best-fit absorption profiles of Ca II K and H. The vertical dashed lines are the central wavelengths of K and H lines. The median impact parameter is displayed in the lower right corner. Note that the scale of y-axis changes as a function of impact parameters.
2.2 Constructing composite spectra
To extract the absorption lines from the quasar spectra, we first employ a dimensional reduction technique, called nonnegative matrix factorization (NMF) (Lee & Seung, 1999), with code [^3] https://github.com/guangtunbenzhu/NonnegMFPy developed by Zhu & Ménard (2013a) and Zhu (2016). With NMF, we normalize the observed quasar spectrum by the reconstructed continuum, derived from a basis set of NMF eigenspectra built from SDSS DR7 quasars (Schneider et al., 2010; Zhu & Ménard, 2013a), to remove the spectral features intrinsic to the quasar. To further remove fluctuations not captured by the NMF method, we obtain a smooth continuum by applying a median filter with a window size of 71 pixels ($\sim 56\,\rm\AA$) to the NMF normalized spectrum and divide the NMF normalized spectrum by the smooth continuum. The window size is empirically chosen to be large enough to ensure sensitivity to the continuum rather than the absorption itself, which spans about 5 pixels (a full width at half maximum of $\sim\mathrm{3.5\,\AA}$). We then shift the normalized quasar spectra to the rest frame of the foreground galaxies, and resample them onto a common wavelength grid with a resolution matching that of the DESI spectra when the doublet line is redshifted to $z=0.2$ ($\sim\mathrm{0.67\,\AA}$). We obtain composite spectra with a median estimator to mitigate the contributions from spectral outliers that occasionally occur. Finally, to account for a sub-percent of zero-point offset, the final composite spectrum is obtained by normalizing the composite spectrum by the median value of the pixels within 3900 and 4100 $\rm\AA$ with the pixels within 5 $\rm\AA$ of Ca II K and H lines being masked. The uncertainties of the median composite spectra are estimated by bootstrapping the samples 500 times.
To measure the Ca II rest equivalent width, we fit the composite spectra with two Gaussian profiles, fixing the central wavelengths at 3934.78 Å (Ca II K) and 3969.59 Å (Ca II H). The amplitudes and line widths are free parameters, but the widths of both lines are enforced to be the same. Throughout this work, we report the sum of the rest equivalent widths of the Ca II K and H lines, denoted as $W_{0}^{\rm K+H}(\mathrm{Ca\ II})$. The uncertainties of the rest equivalent width are based on fitting the 500 bootstrapped samples and taking the standard deviation. Figure 2 shows examples of composite spectra as a function of impact parameter; the best-fit profiles are shown by the blue lines.
Figure 3: Dependence of the Ca II K and H equivalent width on stellar masses as a function of impact parameters. Left: star-forming galaxies. Right: quiescent galaxies. The median stellar masses are listed in the lower left. The blue shaded bands in the right panel correspond to the two highest mass bins for star-forming galaxies in the left panel. The cyan and yellow, square data points are the K line measurements from Rubin et al. (2022) multiplied by a factor of 1.69 to compare with our doublet measurements.
Figure 4: Best-fit amplitudes $A_{100}$ as a function of stellar masses. The data points are plotted with their median stellar masses and respective ranges. The black line is the best-fit power law for the star-forming galaxies.
3 Results
3.1 Dependencies on galaxy properties
With the large number of galaxy-quasar pairs across a wide range of galaxy properties, we first investigate the Ca II rest equivalent width as a function of impact parameter and stellar mass, comparing star-forming and quiescent galaxies. The results are shown in Figure 3 with the left and right panels indicating the Ca II rest equivalent width as a function of impact parameter around star-forming and quiescent galaxies, respectively, with different stellar masses indicated by the colors. The number of pairs is summarized in Table 3.1. We find several trends which are described in the following:
Table 1: Mass bins and best-fit values for star-forming and quiescent galaxies
[…]
$r_{p}$ Star-forming $0.08_{-0.06}^{+0.32}$ $\phantom{0}8.1_{-1.6}^{+0.4}$ $\phantom{0}0.61\pm 1.23$ $-2.15\pm 0.20$ 139946
$\phantom{0}9.0_{-0.5}^{+0.5}$ $\phantom{0}1.76\pm 0.46$ 263731
$\phantom{0}9.9_{-0.4}^{+0.6}$ $\phantom{0}6.79\pm 0.57$ 285305
$10.7_{-0.2}^{+0.8}$ $12.43\pm 1.81$ 47469
$0.06_{-0.04}^{+0.09}$ $\phantom{0}8.1_{-1.6}^{+0.4}$ $\phantom{0}0.28\pm 1.30$ 139874
$\phantom{0}9.0_{-0.5}^{+0.5}$ $\phantom{0}1.78\pm 0.52$ 249580
$\phantom{0}9.8_{-0.3}^{+0.7}$ $\phantom{0}5.70\pm 0.89$ 147123
$10.7_{-0.2}^{+0.8}$ $12.94\pm 5.09$ 12413
$0.21_{-0.06}^{+0.19}$ $\phantom{0}9.5_{-1.0}^{+0.2}$ $\phantom{0}4.93\pm 1.13$ 32145
$10.1_{-0.4}^{+0.4}$ $\phantom{0}8.25\pm 0.95$ 120188
$10.7_{-0.2}^{+0.8}$ $12.49\pm 2.02$ 35056
Quiescent $0.15_{-0.13}^{+0.25}$ $\phantom{0}9.6_{-1.6}^{+0.4}$ $\phantom{0}6.62\pm 2.85$ $-1.86\pm 0.32$ 61976
$10.4_{-0.4}^{+0.3}$ $\phantom{0}4.66\pm 1.65$ 134889
$10.9_{-0.2}^{+1.0}$ $\phantom{0}8.01\pm 2.25$ 116110
$0.09_{-0.07}^{+0.08}$ $10.0_{-2.0}^{+0.5}$ $\phantom{0}4.34\pm 2.83$ 117276
$10.8_{-0.3}^{+1.1}$ $\phantom{0}8.79\pm 3.67$ 61702
$0.24_{-0.07}^{+0.16}$ $10.3_{-1.7}^{+0.2}$ $\phantom{0}6.31\pm 2.78$ 31112
$10.9_{-0.4}^{+1.1}$ $\phantom{0}5.50\pm 1.57$ 102885
AGN $0.07_{-0.05}^{+0.33}$ $10.2_{-3.2}^{+1.3}$ $\phantom{0}4.80\pm 1.27$ $-2.49\pm 1.29$ 39559
$r_{p}/r_{\mathrm{vir}}$ Star-forming $0.20_{-0.18}^{+0.20}$ $\phantom{0}8.8_{-2.3}^{+0.4}$ $\phantom{0}4.05\pm 0.93$ $-2.29\pm 0.24$ 69724
$\phantom{0}9.8_{-0.6}^{+0.4}$ $\phantom{0}6.53\pm 0.61$ 154246
$10.6_{-0.4}^{+0.9}$ $\phantom{0}6.21\pm 0.74$ 143798
Quiescent $0.22_{-0.20}^{+0.18}$ $\phantom{0}9.7_{-1.7}^{+0.3}$ $\phantom{0}7.25\pm 4.48$ $-1.90\pm 0.47$ 28399
$10.5_{-0.5}^{+0.2}$ $\phantom{0}5.13\pm 1.29$ 143912
$11.0_{-0.3}^{+1.0}$ $\phantom{0}1.08\pm 0.62$
Mass dependence: The absorption strengths of Ca II correlate with the masses of star-forming galaxies and quiescent galaxies differently. The $W_{0}^{\rm K+H}$ exhibits an increasing trend with the stellar mass of star-forming galaxies, whereas this trend is much weaker for quiescent galaxies, albeit with large uncertainties. To further quantify these correlations, we fit the $W_{0}^{\rm K+H}$ with a simple power-law relation,
$\langle W_{0}\rangle(r_{p})=A_{100}\left(\frac{r_{p}}{100\ \mathrm{kpc}}\right)^{\alpha},$ (2)
where $A_{100}$ is the equivalent width around galaxies at $r_{p}=100\,\mathrm{kpc}$. We first perform a global fitting to obtain the best-fit $\alpha$ parameter for each type of galaxies, assuming that the radial variations for the same type of galaxies are the same. We then fix the $\alpha$ parameters and obtain the best-fit amplitude for each stellar mass bin. Note that we do not include measurements within 10 kpc as the sight-lines are expected to also intersect the ISM of foreground galaxies, affecting our interpretations for the CGM. The best-fit profiles are shown by the color solid lines in Figure 3 with the best-fit parameter values listed in Table 3.1. The best-fit amplitudes, $A_{100}$, as a function of stellar mass are shown in Figure 4 with the colors representing the corresponding subsamples as in Figure 3. The results demonstrate that Ca II absorption strengths around star-forming galaxies at $\sim 100\,\mathrm{kpc}$ exhibit a mass dependence which can be described by
$A_{100}=C\left(\frac{M_{*}}{10^{10}\,M_{\odot}}\right)^{\beta},$ (3)
with best-fit $\beta$ parameter values being $0.49\pm 0.06$, and $C=6.57\pm 0.47\ \rm m\AA$. On the other hand, the absorption strengths around quiescent galaxies do not depend significantly on stellar mass.
Galaxy type dependence: In addition to the difference of mass dependence, for massive galaxies ($>10^{10}M_{\odot}$) with similar mass, at $r_{p}<30\rm\,kpc$, the absorption strengths around star-forming galaxies are higher than those around quiescent galaxies. By way of illustration, in the right panel of Figure 3, we show the Ca II measurements star-forming galaxies with $10^{10.7}M_{\odot}$ and $10^{9.9}M_{\odot}$. Comparing with the two highest stellar-mass bins of quiescent galaxies, the Ca II absorption within 50 kpc around $10^{10.7}M_{\odot}$ and $10^{9.9}M_{\odot}$ star-forming galaxies is twice as strong as the absorption around $10^{10.9}M_{\odot}$ and $10^{10.4}M_{\odot}$ quiescent galaxies, despite the star-forming galaxies having lower masses. This indicates that the cool gas distributes around star-forming and quiescent galaxies differently when one controls for stellar mass.
SFR dependence: We further examine the dependence of Ca II absorptions on star formation activity by dividing star-forming galaxies into low and high SFR groups using the equation: $\log_{10}{(\mathrm{SFR}/M_{\odot}\,\mathrm{yr^{-1}})}=0.5+0.95\,\left[\log_{10}{(M_{*}/M_{\odot})}-10.0\right].$ The samples for both SFR groups are then divided into three mass bins with matched median masses of $10^{9}$, $10^{9.9}$, and $10^{10.6}\,M_{\odot}$. We follow the previous steps to obtain the Ca II absorption strength as a function of impact parameter, fitting the trend globally with $\alpha=-2.15$. The fitted $A_{100}$ values are shown in Figure 5. By simultaneously fitting $A_{100}$ and SFR across the three mass bins, we find a positive correlation between Ca II absorption strength and SFR, following
$A_{100}=D(M_{*}/M_{\odot})\times\mathrm{SFR}^{0.27\pm 0.03},$ (4)
with $D(10^{9})=2.27\pm 0.24\,\mathrm{m\AA}$, $D(10^{9.9})=5.47\pm 0.25\,\mathrm{m\AA}$ and $D(10^{10.6})=7.72\pm 0.62\,\mathrm{m\AA}$. This result further supports the connection between the cool CGM and the star formation activity of host galaxies.
Figure 5: Best-fit amplitudes $A_{100}$ as a function of SFR for star-forming galaxies. The solid and open points represent the low- and high-SFR groups, respectively, with matched median stellar masses. The lines correspond to the best-fit power law obtained from simultaneous fitting.
AGN dependence: The large DESI dataset enables us to compare the Ca II absorption signal around galaxies with and without an AGN, as identified with the BPT diagnosis (Kewley et al., 2006). The active galaxies sample includes sources in the composite and AGN regions of the BPT diagram, with a median stellar mass of $\sim 10^{10.2}\,M_{\odot}$ and a range from $10^{7}$ to $10^{11.5}\,M_{\odot}$. The results are shown in Figure 6 with the black data points showing the measurements around AGNs and the blue bands showing the measurements around star-forming galaxies with median masses of $10^{9.9}$ and $10^{10.7}\,M_{\odot}$ for comparison. In addition, the dotted line in Figure 6 shows the expected Ca II profile for $10^{10.2}\,M_{\odot}$ star-forming galaxies based on the best-fit relation (Equation 3).
The results show that the Ca II rest equivalent width around AGNs is consistent with that around star-forming galaxies with similar masses, indicating that current AGN activity does not significantly affect the properties of the cool gas traced by Ca II around galaxies. This result is consistent with previous studies, which will be discussed in Section 4.
[…]
Figure 7: Dependence of the Ca II K and H equivalent width on stellar mass as a function of impact parameters, separating the samples into two redshift bins: $z=0.15$ for star-forming galaxies (left), and $z=0.17$ for quiescent galaxies (right). The median stellar masses are listed in the lower left. The solid lines follow the exponents fitted in Figure 3.
Figure 8: Best-fit amplitudes as a function of median stellar masses for two redshift bins. The solid and open points represent low- and high-redshift samples respectively. The solid and dashed lines in the left panel are the best-fit power laws for the low- and high-redshift samples.
Figure 9: Dependence of Ca II on the azimuthal angles for galaxies with $\log_{10}(M_{*}/M_{\odot})\geq 9.5$. Examples of the the galaxy-quasar pairs are shown at the top: from left to right are edge-on galaxies categorized into three azimuthal angles: disk, between disk and pole, pole; followed by face-on galaxies. The galaxies are located at the centre, while the quasars are indicated by the cyan arrows. In the bottom panel, the grey band shows the measurements of the face-on samples.
3.2 Redshift dependence
With the correlations between Ca II absorption and galaxy properties explored, we now investigate the correlation between Ca II absorption and redshift. We separate star-forming and quiescent galaxies at $z=0.15$ and $z=0.17$, respectively, and measure Ca II absorption as a function of stellar mass. The left and right panels of Figure 7 show the results for star-forming and quiescent galaxies, while the upper and lower panels show the results for the low- and high-redshift bins respectively. We also obtain the best-fit amplitudes of power laws (Eq. 2) with fixed $\alpha$ parameters obtained previously for star-forming and quiescent galaxies. The best-fit amplitudes are shown in Figure 8. The solid and open data points indicate the measurements at low and high redshifts respectively.
For star-forming galaxies, the best-fit amplitudes $A_{100}$ of the Ca II absorption strengths increase with stellar masses for both redshift bins. In addition, with a fixed stellar mass, there is a tentative trend showing that the best-fit amplitudes $A_{100}$ at $z>0.15$ are slightly higher than those at $z<0.15$. We quantify this trend with
$A_{100}=C(z)\left(\frac{M_{*}}{10^{10}\,M_{\odot}}\right)^{\beta},$ (5)
and find $\beta=0.45\pm 0.06$. The best-fit values of $C$ are $6.05\pm 0.67\ \mathrm{m\AA}$ ($z<0.15$) and $7.07\pm 0.52\ \mathrm{m\AA}$ ($z>0.15$). This results in
$C(z)=(5.69\pm 0.89)\,\mathrm{m\AA}\,(1+z)^{1.11\pm 0.95}.$ (6)
We note that despite the uncertainty, the redshift trend is similar to the redshift evolution observed in the cool gas around galaxies traced by Mg II absorption lines (e.g., Lan, 2020; Schroetter et al., 2021). For example, Lan (2020) showed that the covering fraction of Mg II absorption lines with $W_{\lambda 2796}>1\,\rm\AA$ around star-forming galaxies evolve with redshift as $(1+z)^{2.2\pm 0.4}$. On the other hand, we do not observe any redshift dependence for Ca II around quiescent galaxies which is inconsistent with the results of Lan (2020) for quiescent galaxies. This might indicate that the cool gas around passive and star-forming galaxies evolve differently. However, further confirmation of such a scenario requires higher S/N measurements of Ca II absorption.
In Lan (2020), the author proposed that the redshift evolution of Mg II covering fraction is connected with the redshift evolution of star-formation rate of galaxies (e.g., Whitaker et al., 2012; Scoville et al., 2017; Koprowski et al., 2024) given that they have similar redshift evolution. For our Ca II measurements, we also estimate the SFR of star-forming galaxies at two redshift bins with stellar mass $>10^{10}M_{\odot}$ and find that the ratio between the median SFR, $\mathrm{SFR}(z>0.15)/\mathrm{SFR}(z<0.15)$, is $\sim 1.8$. Together with our results indicating a SFR and absorption strength correlation with $\propto\rm SFR^{0.27}$ (Fig. 5), the absorption strength difference ratio is expected to be $\sim 1.8^{0.27}\sim 1.17$, which is consistent with the redshift trend observed in Figure 8. This consistency suggests that the redshift evolution of Ca II absorption around star-forming galaxies might be closely related to the star-formation activities of galaxies.
3.3 Azimuthal angle dependence
One can further investigate possible mechanisms contributing to gas observed in the CGM by measuring gas absorption around star-forming galaxies as a function of azimuthal angle. This is motivated by the fact that galaxy simulations predict that galactic outflows driven by supernova and black-hole feedback preferentially escape galaxies along the minor axis (e.g., Nelson et al., 2019). To do so, we examine the azimuthal dependence of our sample by choosing star-forming galaxies with $\log_{10}\left({M_{*}/M_{\odot}}\right)\geq 9.5$ and with inclination angle larger than $40^{\circ}$. The inclination angles of galaxies and azimuthal angles between galaxy-quasar pairs are calculated based on the shape information measured via Tractor [^4] https://www.legacysurvey.org/dr9/catalogs/ (Lang et al., 2016) applied to the DESI Legacy Imaging Surveys (Dey et al., 2019). In general, the Tractor performs model-based photometry by simultaneously fitting sources across multiple bands and classifies them as point or extended sources with the point spread function taken into account. Extended sources are further modeled using exponential, de Vaucouleurs, or composite profiles to best represent their light distribution. We then divide the selected sample into three azimuthal angle bins: $|\phi|\in[0^{\circ},30^{\circ}]$, $[30^{\circ},60^{\circ}]$ and $[60^{\circ},90^{\circ}]$, which correspond to the disk, between disk and pole, and the pole of the foreground galaxies. Galaxy image examples can be found in the upper panel of Figure 9, where the foreground galaxies are located at the centers, while the background quasars are pointed out by cyan arrows.
The results are shown in the lower panel of Figure 9 with colors indicating measurements for different azimuthal angles and the grey band for the measurements of the face-on galaxies. While there is a mild trend showing that the absorption along the major axis decreases faster than the absorption along the minor axis toward larger impact parameters, there is no significant correlation between Ca II absorption and the azimuthal angle. This result is inconsistent with previous results of the cool CGM obtained at higher redshifts, showing preferentially stronger Mg II absorption along the minor axis than along the major axis within 50 kpc (e.g., Bordoloi et al., 2011; Lan et al., 2014; Lan & Mo, 2018). In addition, our results are inconsistent with Ca II measurements from Zhu & Ménard (2013b). The authors reported that the absorption strength of Ca II is stronger along the minor axis than the major one. However, we note that the conclusion in Zhu & Ménard (2013b) is driven by a single measurement at $r_{p}\sim 15\ \mathrm{kpc}$, a trend that is not observed in our measurements with more pairs. We will further discuss the implications of our azimuthal angle measurements in Section 4.2.
3.4 Gas distribution with respect to the size of halos
Given that the same impact parameter in the physical space corresponds to different regions in the halos of galaxies with different masses, we measure the Ca II absorption as a function of impact parameter normalized by the virial radius, $r_{p}/r_{\mathrm{vir}}$, of the dark matter halos of the central galaxies. To estimate the virial radius, we follow Chang et al. (2025) and derive halo masses using Bayesian inference. Briefly, the halo masses are estimated from stellar masses combining information of a prior and a likelihood function, which are constructed using the halo mass function from Tinker et al. (2010) and the stellar-to-halo mass relation from Moster et al. (2013), respectively. For further details, we refer the readers to Chang et al. (2025). Using the estimated halo masses, we determine the virial radius with the analytical formula from Bryan & Norman (1998).
The results are shown in Figure 10. The Ca II measurements around star-forming and quiescent galaxies are shown in the left and right panels respectively, fitted with
$\langle W_{0}\rangle\left(\frac{r_{p}}{r_{\mathrm{vir}}}\right)=A_{0.5\rm vir}\left(\frac{r_{p}}{0.5r_{\mathrm{vir}}}\right)^{\eta}.$ (7)
We find that after the normalization, Ca II absorption profiles around star-forming galaxies with different stellar masses become consistent with each other, indicating that the cool gas distributes similarly in the halos, and the behavior of the cool CGM is closely tied to the halo masses of the host galaxies. On the other hand, for quiescent galaxies, there is a different trend, showing that the absorption strengths decrease with stellar masses, especially in the inner regions of the halos $<0.3\,r_{\mathrm{vir}}$. In addition, again, for galaxies with $>10^{10}\,M_{\odot}$, the Ca II absorption around star-forming galaxies is approximately 1.5-2 times stronger than that around quiescent galaxies as observed in the physical space. These results demonstrate that while cool gas traced by Ca II exists in both types of galaxies, they behave differently around different types of galaxies.
Our Ca II results are consistent with results obtained from Mg II absorption lines, showing that the Mg II absorption profiles around star-forming galaxies with different masses are similar when normalized by the distance $r_{p}/r_{\mathrm{vir}}$, while the absorption decreases with the mass of galaxies for quiescent ones (e.g., Churchill et al., 2013a, b; Lan, 2020; Anand et al., 2021).
Figure 10: Dependence of the Ca II K and H equivalent width on stellar masses as a function of normalized distances. Left: star-forming galaxies. Right: quiescent galaxies. The median stellar masses are listed in the lower left.
3.5 Ca II mass in the halos
We now estimate the Ca II mass in the halos. To this end, we first calculate the Ca II column density based on the Ca II K line rest equivalent width as $\langle W_{0}^{\mathrm{K}}\rangle=\frac{2}{3}\times\langle W_{0}^{\mathrm{K+H}}\rangle$ using the following equation adopted from (Draine, 2011)
$N_{\mathrm{Ca\,II}}=1.13\times 10^{12}\ \mathrm{cm^{-1}}\ \frac{W_{0}^{\mathrm{K}}}{f\lambda^{2}},$ (8)
where $f$ is the oscillator strength of Ca II K being 0.648, and $W_{0}^{\mathrm{K}}$ and $\lambda$ are in units of $\rm\AA$. This relation works for unsaturated, optically thin absorption lines and provides a lower limit of column density when the lines are saturated. Our measurements are in the unsaturated region with the strongest $\langle W_{0}^{\mathrm{K}}\rangle$ being $0.2\,\rm\AA$ at $r_{p}>10\,\mathrm{kpc}$ (e.g., Zhu & Ménard, 2013b). We then adopt the best-fit profiles as shown in Figure 10, and integrate the Ca II mass from $0.1r_{\mathrm{vir}}$ to $1r_{\mathrm{vir}}$:
$\langle M_{\mathrm{Ca\,II}}\rangle=2\pi m_{\mathrm{Ca}}\int_{0.1r_{\mathrm{vir}}}^{r_{\mathrm{vir}}}N_{\mathrm{CaII}}\left(r_{p}\right)r_{p}\mathrm{d}r_{p},$ (9)
where $m_{\mathrm{Ca}}$ is the atomic mass of Ca. The results are shown in Figure 11. We emphasize that given that Ca II only contains a fraction of Ca in the ionization state, the estimated Ca II mass represents the lower limit of the total Ca mass.
In addition, we estimate the total Ca mass in the ISM of star-forming galaxies by combining the stellar mass-metallicity relation and atomic gas sequence from Scholte et al. (2024) with the solar Ca-to-O abundance ratio of 0.004 (Lodders, 2019). We also apply a dust depletion correction to the metallicity using the relation from Peimbert & Peimbert (2010). The estimated total Ca mass in the ISM of star-forming galaxies is shown by the grey line. For the CGM, we also account for dust depletion in our Ca II measurements of the halos by applying a depletion factor. Using Cloudy simulations (Chatzikos et al., 2023), we determine a depletion factor of $\sim 89$ by comparing the intrinsic Ca II-to-Zn II ratio in the CGM with the observed ratio from Lan & Fukugita (2017). A detailed description of the simulations is provided in Appendix A. After the correction, the amount of the Ca II mass in the halos is comparable to the amount of Ca mass in the ISM. Moreover, the relationship between the Ca II mass and stellar mass of star-forming galaxies follows the relationship of the ISM metal mass and stellar mass of star-forming galaxies. This result demonstrates that the metal enrichment of the ISM and CGM is closely linked together, reflecting the cumulative effect of feedback mechanisms.
Figure 11: Ca II mass in CGM. The solid data points are estimated Ca II mass around star-forming galaxies (blue) and quiescent galaxies (red) and the open data points are values corrected for dust depletion. The grey line is the Ca mass in ISM inferred from the mass-metallicity relation and atomic mass sequence.
4 Discussion
4.1 Comparisons with previous results
By combining DESI spectra of background quasars, we have measured the Ca II absorption strengths around galaxies as a function of stellar mass, galaxy types and redshift, and revealed correlations between Ca II absorption strengths and galaxy properties. In the following, we compare our results with similar measurements from previous studies and discuss the implications.
Cool CGM traced by Ca II at $z<0.2$: Our observed trends are consistent with the trends reported in the previous Ca II results at $z\sim 0.1$ using the SDSS dataset by Zhu & Ménard (2013b). For example, Zhu & Ménard (2013b) separated galaxies into two mass bins and found that galaxies with $10^{10.5}$ and $10^{9.8}\,M_{\odot}$ have best-fit $A_{100}$ parameter values of $15\,\rm m\AA$ and $7\,\rm m\AA$ at 100 kpc. This yields $\beta\sim\rm 0.47$, which is consistent with our best-fit $\beta$ parameter value. We note that while in their analysis, only a mass cut is applied, we expect that the trend in their measurements is mostly driven by signals from star-forming population which has a larger pair number. In addition, they reported that Ca II around star-forming galaxies is approximately two times stronger than Ca II around quiescent galaxies, a trend that we have found in DESI data. We note that limited by the SDSS dataset, Zhu & Ménard (2013b) explored either the dependences between Ca II absorption and mass or SFR, while the DESI dataset enables us to perform the exploration as a function of mass and SFR independently across a wider parameter space than previously explored. We also note that the $\alpha$ parameter of our best-fit profiles is $\sim-2$ for both types of galaxies while Zhu & Ménard (2013b) obtained $\sim-1.4$. The difference is mostly due to the fact that we exclude the $<10\ \,\rm kpc$ measurements during the fitting, while Zhu & Ménard (2013b) included the measurements in their fitting. In Appendix B, we directly compare DESI measurements with SDSS measurements from Zhu & Ménard (2013b), and show that the measurements from the two datasets are consistent.
The only inconsistency between our results and the results in Zhu & Ménard (2013b) is the azimuthal angle dependence. Given that the trend observed in Zhu & Ménard (2013b) is driven by a single measurement in the innermost bin, it is possible that such a trend is affected by small statistics limited by the SDSS data. The larger DESI sample is expected to be less sensitive to the effect. Our measurements are also consistent with the Ca II measurements within 10 kpc of galaxies obtained by Rubin et al. (2022) [^5] We multiply the Ca II K measurements in Rubin et al. (2022) by by a factor of 1.69, which is the weighted mean of the ratio of H+K to K lines from the innermost bins of our measurements. as shown in Figure 3, as well as the Ca II measurements in our Milky Way ISM with typical $W_{0}^{\rm H+K}\sim 0.8-1\rm\,\AA$ from Murga et al. (2015) [^6] Given that the Milky Way sightlines intercept approximately half of the ISM, we multiply the Ca II measurements by two..
Cool CGM traced by Mg II at $z\gtrsim 0.4$: Explorations of the galaxy-cool CGM connections have been conducted with Mg II absorption lines as the tracers of the cool CGM (e.g., Bordoloi et al., 2011; Kacprzak et al., 2012; Rubin et al., 2014; Nielsen et al., 2016; Dutta et al., 2020; Huang et al., 2021), and have revealed similar correlations as found in this work:
  • Mass dependence: Studies have shown that the covering fraction and the absorption profiles of the cool CGM traced by Mg II absorbers increase with stellar mass (and halo mass) of systems. For example, using $\sim 100$ galaxy-quasar pairs, Chen et al. (2010) showed that the characteristic radius of gas traced by Mg II scales with galaxy B-band luminosities as $\propto L_{B}^{0.35}$. With $\sim 200$ galaxy-quasar pairs obtained with MUSE observations, Dutta et al. (2020) also concluded that the stellar mass is the primary factor affecting Mg II absorption around galaxies. Similarly, via a statistical cross-correlation analysis, Lan (2020) demonstrated that the covering fraction of Mg II absorbers measured in physical space correlates with the galaxy mass with $\propto M_{*}^{0.5}$ for star-forming galaxies. Finally, based on cool gas measurements around galaxies and quasars living in halos with different masses, Prochaska et al. (2014) also argued that the halo mass is the primary factor affecting the cool gas distribution, a conclusion which is also reached by Churchill et al. (2013a).
    Together with the Ca II measurements, this mass dependence of the cool CGM has been observed across a wide range of redshifts from $z\sim 0$ up to $z\sim 2$, covering at least 10 Gyr cosmic time. This provides hints on the underlying mechanisms which are expected to operate from $z\sim 2$ and persist through cosmic time.
  • Galaxy type dependence: In addition to the mass dependence, the correlation between galaxy types and the cool CGM, especially in the inner regions, has been reported. Several studies show that the average Mg II absorption strength in the inner CGM around star-forming galaxies is stronger than that around quiescent galaxies (e.g., Bordoloi et al., 2011; Lan et al., 2014; Lan, 2020; Rubin et al., 2018; Anand et al., 2021; Huang et al., 2021). A difference of a factor of $\sim 4$ is observed within 50 kpc, and the value is larger with smaller impact parameters. We note that such a trend is found in Ca II measurements as shown in Section 3. However, the differences in Ca II absorption between star-forming and quiescent galaxies are at most a factor of 2 at $r_{p}\sim 20$ kpc, which is smaller than the results from MgII absorption at higher redshifts. This might suggest that the dependence of the cool CGM on galaxy types becomes weaker at lower redshifts, possibly associated with the declines of the SFR of galaxies and the contribution of galactic outflows. We discuss this implication in the following subsection.
4.2 Possible redshift evolution of the azimuthal angle dependence
Blue-shifted absorption lines, the observed signature of galactic outflows, have been detected ubiquitously in the spectra of star-forming galaxies, via the so-called down-the-barrel observations (e.g., Weiner et al., 2009; Rubin et al., 2014; Zhu et al., 2015). With the typical observed velocity ($\sim 300$ km/s; e.g., Rubin et al. 2014), gas associated with outflows is expected to travel into the halos and possibly reaches to 50 kpc within a few hundred Myr (e.g., Lan & Mo, 2019). The azimuthal angle dependence of gas distribution has been considered as one of the possible imprints of galactic outflows in the CGM. Such a distribution has been observed in the cool gas traced by Mg II absorption lines. For instance, Bordoloi et al. (2011) measured the Mg II absorption along the minor and major axes of star-forming galaxies at $z\sim 0.7$ with HST ACS images and found that within 50 kpc, the Mg II absorption is stronger along the minor axis than the major axis. The rest equivalent width of Mg II is $\sim 3$ times higher along the minor axis than the major axis at $\sim 25$ kpc. A similar trend and the ratio between Mg II rest equivalent width along the minor and major axes are also found in Lan & Mo (2018) at $z\sim 1$. With individual detections, the bimodal azimuthal dependence of Mg II absorption is reported at $z\sim 0.2$ (e.g., Bouché et al., 2012; Kacprzak et al., 2012). However, the ratio of the Mg II rest equivalent width along the minor to major axes is not as large as the value observed at $z>0.7$ (e.g., Martin et al., 2019; Huang et al., 2021). The azimuthal dependence of Ca II measurements reported in Section 3.3 is more consistent with the $z\sim 0.2$ Mg II results.
Here we propose that the different degree of the azimuthal angle dependence is possibly due to the redshift evolution of galactic outflows driven by star-formation activities. As discussed above, the azimuthal angle dependence is observed mostly at $z>0.7$ (Bordoloi et al., 2011; Lan & Mo, 2018), while the low-redshift CGM has weak or no detectable azimuthal angle dependence as shown in this work and previous studies (e.g., Huang et al., 2021; Borthakur et al., 2015; Pointon et al., 2019). This can be due to the fact that for star-forming galaxies with similar mass, the average SFR at $z\sim 0.8$ is $\sim 3$ times higher than that at $z\sim 0.2$ (e.g., Moustakas et al., 2013). Therefore, three times more gas mass associated with outflows is expected to contribute to the CGM preferentially allocated along the minor axis. Moreover, the velocities of galactic outflows evolve with redshifts. With a fixed SFR, the maximum outflow velocity at $z\sim 1$ is approximately 1.5-2 times faster than the outflow velocity at $z\sim 0$ (e.g., Sugahara et al., 2019). Taking these two factors together, at higher redshifts, galactic outflows are expected to distribute more gas mass into the CGM at larger distances. In other words, the absence of azimuthal angle dependence observed in Ca II at low redshifts indicates that gas associated with galactic outflows does not contribute significantly in the inner CGM. This lack of galactic outflow contribution can also explain why Ca II absorption in the inner CGM around star-forming and quiescent galaxies at $z\sim 0.2$ does not have such a large difference observed at $z>0.7$ with Mg II absorption.
This suppression of galactic outflows towards low redshifts is aligned with the results from the FIRE (Feedback In Realistic Environments) simulations (Hopkins et al., 2014). Muratov et al. (2015) and Anglés-Alcázar et al. (2017) found that high-redshift galaxies drive strong galactic outflows produced by bursty star formation, while at low redshifts, massive galaxies shift into steady star-formation that does not drive strong outflows. Stern et al. (2021) proposed that this transition of the star-formation mode is linked to the virialization of the inner CGM, which occurs at lower redshifts, and after the virialization of the outer CGM. Before the virialization of the inner CGM, the inner CGM is dominated by clumpy cool gas, making galactic outflows relatively easy to travel through. After the virialization of the inner CGM, the hot gas around galaxies confines the galactic outflows. This transition of the inner CGM virialization for $L^{*}$ galaxies occurs at $z\sim 0.8-1$ in the FIRE simulations which coincides with the epoch separating the behaviors of cool gas around star-forming galaxies. While further testing and validation of this outside-in CGM virialization scenario requires direct comparisons between the predicted and observed properties of cool gas tracers, such as Mg II and Ca II, around galaxies across time and mass, the current observed properties of the cool CGM are consistent with the results of the FIRE simulations qualitatively.
4.3 Impact of AGNs on the properties of the CGM
Our results show that there is no significant difference in the Ca II absorption between AGNs and star-forming galaxies. This indicates that current AGN activities do not have strong impact on the properties of the cool CGM. This result is consistent with previous measurements of gas around galaxies with AGN activities. For example, Berg et al. (2018) selected 19 AGNs via BPT diagnosis and measured the CGM properties with COS UV spectroscopy, finding that the absorption line properties of AGNs are consistent with those of the control sample within 160 kpc. Similarly, Kacprzak et al. (2015) estimated the covering fraction of Mg II absorption around 14 AGNs, and obtained a value slightly lower than that of field galaxies, though within the uncertainties. For luminous AGNs, i.e. quasars, the cool gas distribution has also been measured via quasar-quasar pairs. Previous results show that the properties of cool gas around quasars are similar to those around galaxies with similar stellar mass, suggesting that the halo mass is the primary factor driving the properties of the cool gas (e.g., Farina et al., 2014; Prochaska et al., 2014). On the other hand, Johnson et al. (2015) reported a correlation between the covering fraction of Mg II absorption and quasar luminosity, and argued that quasar feedback can be one of the mechanisms driving the correlation. While validating such a physical connection requires further investigations, it suggests that energy from powerful / luminous quasars may be needed to sufficiently leave imprints on the properties of the cool CGM. Finally, considering radio-mode feedback, Chang et al. (2024) measured the properties of the cool gas traced by Mg II around radio galaxies and the control sample, and found no detectable difference between the properties of the gas around the two populations, illustrating that recent radio AGN activities do not impact the properties of the cool gas strongly. These measurements again support the picture that the properties of galaxies, especially the stellar mass / halo mass and SFR, are the dominant factors correlating with the properties of the cool gas.
5 Conclusions
Utilizing the DESI Y1 galaxy and quasar spectroscopic datasets, we constructed high S/N composite spectra of quasars in the background of galaxies from which we detected and measured the absorption strengths of Ca II, tracers of cool gas, around galaxies. We explored the relationships between the properties of the cool CGM and the properties of galaxies, including stellar mass, SFR, redshift, and AGN activities, and compared our measurements at $z<0.4$ with similar measurements obtained at higher redshifts. Our results are summarized in the following:
  1. 1.
    We find that the Ca II rest equivalent widths decrease with impact parameters, consistent with the cool CGM measurements from previous studies.
  2. 2.
    For star-forming galaxies, the absorption strengths correlate with the stellar mass of galaxies from $10^{8}\,M_{\odot}$ to $10^{11}\,M_{\odot}$ with $\propto M_{*}^{0.5}$, while such a trend is weaker for quiescent galaxies. With a fixed stellar mass, we also find a positive correlation between absorption strengths and SFR for star-forming galaxies, following $\propto\mathrm{SFR}^{0.3}$. In addition, for $>10^{10}\,M_{\odot}$ galaxies, the absorption strengths are stronger around star-forming galaxies than those around quiescent galaxies within $30\,\mathrm{kpc}$. The correlations between Ca II and galaxy properties are detected at $z\sim 0.1$ and $z\sim 0.3$.
  3. 3.
    We find that there is no azimuthal angle dependence of absorption strength around star-forming galaxies, which is inconsistent with previous measurements at higher redshifts. We argue that this might be due to the redshift evolution of galactic outflows which is expected to carry less mass with slower velocity at lower redshifts.
  4. 4.
    We measure Ca II absorption around AGNs, selected with the BPT diagnosis, and find that with similar galaxy masses, Ca II absorption strengths around AGNs and star-forming galaxies are consistent. This result suggests that current AGN activities do not have strong impacts on the properties of the cool CGM.
  5. 5.
    After normalizing the impact parameter by the virial radius, the absorption profiles around galaxies with different stellar masses from $10^{8}\,M_{\odot}$ to $10^{11}\,M_{\odot}$ become consistent with each other, indicating that the cool CGM distributes similarly in the halos and is closely linked to halo masses. For quiescent galaxies with masses $>10^{11}\,M_{\odot}$, the Ca II absorption in the inner CGM appears to be weaker. This halo mass trend is also observed at higher redshifts, suggesting that the underlying mechanism(s) regulating such a property persists across cosmic time.
  6. 6.
    We estimate the mass of Ca II as a function of stellar mass, and find that after the dust depletion correction, the amount of Ca II in the CGM is comparable with the Ca mass in the ISM of galaxies. Moreover, the cool CGM metal mass and stellar mass relation of star-forming galaxies follows the ISM metal mass and stellar mass relation.
Our results provide a comprehensive characterization of galaxy-cool CGM connections below $z<0.4$ down to $10^{8}\,M_{\odot}$ galaxies, offering novel constraints on the models of galaxy evolution. Together with similar measurements at higher redshifts, the relationships between the properties of galaxies and the cool CGM have been mapped across $\sim 10$ Gyr from redshift 0 to redshift $\sim 2$. Understanding the physical mechanisms behind these relationships over cosmic time will be essential for advancing the knowledge of galaxy formation and evolution.
The analysis adopted in this work can be directly applied to upcoming datasets. For example, at the local Universe, the complete 5-year DESI dataset is expected to further enhance the S/N of the similar measurements by at least a factor of two. In addition, one can combine the images provided by Euclid (Euclid Collaboration et al., 2024) to obtain more precise measurements of gas distribution as a function of azimuthal angle, a key constraint on how gas flows around galaxies. Similar measurements will also be obtained at even higher redshifts. The Prime Focus Spectrograph (PFS) survey (Takada et al., 2014) and the DESI-II (Schlegel et al., 2022), for example, will observe millions of galaxies at $z>2$, enabling explorations of galaxy-CGM connections with large datasets at the peak of the star-formation density of the Universe (e.g., Madau & Dickinson, 2014). Similar to the properties of galaxies, such as stellar mass functions and luminosity functions at different redshifts (e.g., Kelvin et al., 2014; Driver et al., 2022; Weaver et al., 2023), precise measurements of gas around galaxies at different redshifts will serve as fundamental constraints, leading to a more comprehensive understanding of the cosmic baryon cycle.
Data Availability
All data points shown in the figures are available at Zenodo doi: 10.5281/zenodo.14970325.
We thank Siwei Zou and Abhijeet Anand for their constructive comments that greatly helped improve the paper during the DESI Collaboration Wide Review. YVN, TWL, and YLC acknowledge supports from the National Science and Technology Council (MOST 111-2112-M-002-015-MY3, NSTC 113-2112-M-002-028-MY3), the Yushan Fellow Program by the Ministry of Education (MOE) (NTU-110VV007, NTU- [^1] V1007-2, NTU-112V1007-3, NTU-113V1007-4), National Taiwan University research grant (NTU-CC- [^1] L894806, NTU-CC-112L893606, NTU-CC-113L891806, NTU- [^1] L7318, NTU-112L7302). This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. The DESI Legacy Imaging Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS), the Beijing-Arizona Sky Survey (BASS), and the Mayall z-band Legacy Survey (MzLS). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory. Legacy Surveys also uses data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. Legacy Surveys was supported by: the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility; the U.S. National Science Foundation, Division of Astronomical Sciences; the National Astronomical Observatories of China, the Chinese Academy of Sciences and the Chinese National Natural Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. The complete acknowledgments can be found at https://www.legacysurvey.org/. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies. The authors are honored to be permitted to conduct scientific research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation.
Appendix A Cloudy simulation
We perform Cloudy simulations (Chatzikos et al., 2023) to evaluate the dust depletion factor in the CGM. The simulations are initialized with the Haardt & Madau (2012) extragalactic radiation background at $z=0.2$, and run over a hydrogen volume density range of $10^{-2}$ to $10^{0.5}\,\mathrm{cm^{-3}}$ until the neutral hydrogen column density reaches $10^{19}$, $10^{19.3}$ and $10^{19.5}\,\mathrm{cm^{-2}}$. In order to determine the depletion factor, we run two sets of simulations. First, we run a simulation by adopting the default solar composition in Cloudy (Grevesse & Sauval, 1998; Allende Prieto et al., 2001; Holweger, 2001; Allende Prieto et al., 2002) to obtain the intrinsic column density ratio of Ca II and Zn II (the left panel of Figure 12). Next, we reduce the calcium abundance to simulate dust depletion, adjusting it to match the observed average Ca II-to-Zn II column density ratio of $\sim 0.86$ from Lan & Fukugita (2017) (the middle panel of Figure 12). By comparing the intrinsic ratio with the depleted simulations, we derive an average Ca depletion factor of $\sim 89$ and adopt this value to correct the depletion effect.
Figure 12: Cloudy simulation results across a total hydrogen volume density range of $10^{-2}$ to $10^{0.5}\,\mathrm{cm^{-3}}$ for $\log_{10}{(N_{\mathrm{HI}}/\mathrm{cm^{-2}})}=19.0,19.3,19.5$. The left panel shows the intrinsic Ca II-to-Zn II column density ratio, assuming an initial solar abundance. The middle panel shows the observed ratio from Lan & Fukugita (2017) (dashed line) and simulated ratios by reducing the calcium abundance by a factor of $10^{-1.95}$. The right panel presents the depletion factor, obtained by dividing the intrinsic ratio from the left panel by the simulated depleted values from the middle panel.
Appendix B Comparison with SDSS
Here we compare our DESI measurements with the SDSS measurements from Zhu & Ménard (2013b). To match the SDSS galaxy population, we select DESI galaxies with $r<17.7$ (Strauss et al., 2002). As shown in Figure 13, the DESI measurements are consistent with the SDSS measurements. We note that while DESI provides more galaxy-quasar pairs, the typical S/N of individual DESI quasar spectra is lower than the S/N of SDSS quasar spectra. Together, these two factors yield similar S/N values of the SDSS and DESI composite spectra. Nevertheless, the DESI dataset enables us to probe the Ca II absorption around galaxies with lower mass and higher redshifts.
Figure 13: Comparison of Ca II absorption (K line) around $r<17.7$ galaxies with SDSS and DESI data. The grey data points show the measurements from Zhu & Ménard (2013b) and the blue data points show the measurements obtained in this work with DESI data.